# Autoregressive Image Generation Guided by Chains of Thought

### v1

@@ 4.2 Ablation Study
4.2.1 Robustness of IGTR
To evaluate the robustness of proposed IGTR, we conduct experiments using different types of reasoning prompts. Specifically, we use two types of prompts mentioned in Sec. 3.2.3: class reasoning prompt and universal reasoning prompt, as the intermediate reasoning process for our image generation. We evaluate the effectiveness and robustness of IGTR by comparing the results with consistent reasoning prompts (i.e., the same reasoning prompt used during both training and inference) and inconsistent reasoning prompts (i.e., different reasoning prompts used during training and inference). In addition, we also combine two different prompts for training and inference, with each having a 50% probability of being selected. We train the model with  [^1] M parameters for 300 epochs, which is the same settings as those in LlamaGen [35]. The baseline is LlamaGen [35], an AR image generation model based on the Llama architecture without our reasoning prompt.
@@ 4.2.1 Robustness of IGTR
Non Reasoning Prompt 8.69 124.44 0.79 0.47
Reasoning After Generation 8.70 122.79 0.79 0.47
@@
IGTR 6.07 174.48 0.82 0.44
@@
Table 4: Ablation study for differenct CFG settings.
The results are shown in Table 2, with the following observations: (1) Regardless of whether we use consistent or inconsistent reasoning prompts for training and inference, and regardless of which reasoning prompt we choose for the intermediate reasoning process, our method consistently outperforms baseline without reasoning prompt, achieving an approximate 20% improvement in FID scores. This demonstrates that our method exhibits good robustness. (2) The performance of the class reasoning prompt is better than that of the universal reasoning prompt. This is because focusing on the overall distribution of a specific class is more informative than considering the overall distribution of the entire image set, as it more accurately reflects the distribution of the specific class. (3) When the reasoning prompt is consistent between training and inference, the performance is better than when they are inconsistent. This is expected, as consistency between training and inference ensures that the model can better align the reasoning process across both phases, leading to more coherent and stable results. (4) The inferior performance of the mixed prompt is due to the introduction of information conflict and inconsistency. When using two different prompts, the model needs to balance both, which could lead to confusion in the generated content. Compared to a single prompt, the mixed prompt increases the complexity of reasoning, and the model may not have sufficient capacity to effectively integrate the information from both prompts, thus affecting performance.
@@ 4.2.2 Different Number of Prompt Tokens
We conduct an ablation study on the length of the intermediate reasoning prompt, using the universal reasoning prompt as the intermediate reasoning prompt, which samples a set of indices from the image tokenizer’s codebook based on a uniform distribution. To save computational resources, all experiments train the model with  [^1] M parameters for 50 epochs, as 50 epochs are sufficient to reflect the performance of different models. The results are shown in Fig. 4. A prompt length of 0 corresponds to the baseline, which represent the baseline without intermediate reasoning prompt. We can observe that when the reasoning prompt is short (8 tokens), ability of IGTR to handle complex image generation task is weaker due to containing less information, resulting in performance similar to the baseline. Furthermore, as the length of the prompt increases, the FID performance improves gradually. However, the performance gains are not uncapped, which could be found that when the prompt length reaches near 256 the gains begin to slow down. This phenomenon is reasonable as the gains from using distributed information will reach an upper limit. Therefore, in subsequent experiments, considering both performance and efficiency, the length of reasoning prompt is set to 256.
@@
IGTR-B (CFG=1.75)  [^1] M 4.39 222.80 0.85 0.46
IGTR-L (CFG=1.75) 343M 3.23 279.07 0.84 0.52
IGTR-XL (CFG=1.75) 775M 2.85 289.75 0.83 0.54
IGTR-XXL (CFG=1.75) 1.4B 2.76 275.80 0.82 0.55
@@
Figure 5: Sample images randomly generated by IGTR, trained on ImageNet.
Figure 6: Scaling behavior of IGTR. (a) Training loss of models with different sizes. (b) FID scores of models with different sizes during inference.
Figure 7: Visualization of samples generated by IGTR based on different model sizes.
4.2.3 Order of Reasoning and Generation
A potential benefit of IGTR is that it allows the model to better access relevant knowledge acquired during the pretraining process. To investigate this, we test an alternative configuration where the intermediate reasoning prompt is given only after the image generation, isolating whether the model actually depends on the prior reasoning prompt to generate the image. The experiments train the model for 50 epochs. As shown in the ‘Reasoning After Generation’ row in Table 3, we find that the performance of this variant is comparable to the baseline, suggesting that the sequential reasoning itself is valuable in the generation process of the model, rather than merely activating known knowledge.
@@ 4.2.4 Blank Prompt
Another way to think about our IGTR is that the proposed image-related reasoning prompt enables the model to spend more computational resources (i.e., intermediate tokens). To isolate the impact of variable computation from the reasoning process itself, we conduct a variant where we replace the intermediate reasoning prompt with multiple blank tokens of equal length, which are tokens extracted from a completely black image with no distributional information. The experiments train the model for 50 epochs. The experimental result, as shown in ‘Blank Prompt’ of Table 3, indicates that this variant performs similarly to the baseline without our reasoning prompt. This demonstrate that variable computation itself is not the reason for the success of IGTR, and that IGTR expresses intermediate reasoning steps with distributional information to be effective.
@@ 4.2.5 Effect of Classifier-Free Guidance (CFG)
We conduct ablation experiments under different Classifier-Free Guidance (CFG) settings based on random reasoning prompt, as shown in Table 4. It can be observed that as CFG increases, the model’s performance gradually improves, with the best FID achieved when CFG is set to 1.75. However, further increasing CFG leads to a deterioration in FID. Additionally, the increase in CFG results in a trade-off between diversity and fidelity, with higher accuracy and lower recall. Therefore, in subsequent experiments, we set the CFG to 1.75.
@@ 4.3 Main Results
In Table 5, we compare with popular image generation models, including GAN [2, 13, 30], Diffusion models [6, 12, 27, 24], masked AR [4, 18, 19] and AR models [8, 42, 17, 35]. All our experiments are based on the universal reasoning prompt, as it is more general and easier to obtain compared to the class reasoning prompt.
The results show that our models outperform all previous AR methods at different levels of model parameters. For example, compared to Llama-B, Llama-L Llama-XL and Llama-XL with the same parameters, IGTR-B, IGTR-L IGTR-XL and IGTR-XL improves FID by 20%, 15%, 16% and 11%, and IS by 15%, 12%, 28% and 9%, respectively. This indicates that our IGTR of introducing reasoning prompts effectively helps the AR model in learning and generation by enhancing contextual reasoning of generation and improving stability during the generation process through increased inference steps . It is important to note that since LlamaGen does not release the results at 300 epoches for the L and XL models, for a fair comparison, we only present the results for our method at 50 epochs on the XL and XXL models, which have not yet reached the performance limit of our method, leaving room for further improvement.
@@
As shown in Fig. 5, we visualize some sample images randomly generated by IGTR, trained on ImageNet. This demonstrates that IGTR is capable of generating high-quality samples with both high fidelity and diversity. More visualizations are provided in the appendix.
@@
We investigate the scaling behavior of IGTR. As shown in Fig. 6, we show the training loss and FID score based on models of different sizes. It can be observed that IGTR demonstrates good scalability, with lower training loss and better FID scores as the model size increases. This is because we did not modify the formula or structure of the AR model itself, preserving the integrity of the AR framework, and therefore inheriting the scalability of AR methods. Additionally, we visualize the generation results of the same class under different model sizes, as shown in Fig. 7. It can be observed that as the model size increases, the generated results for the same class become more realistic and detailed, with finer textures and better overall consistency.
@@ 5 Conclusion
In this work, inspired by the CoT reasoning technique from NLP of AR model, we propose Image Generation with Thoughtful Reasoning (IGTR), to enhance AR image generation. IGTR improves both the accuracy and stability of image generation without changing the model structure or the autoregressive generation order. By adding specialized image-related reasoning prompts, IGTR enhance contextual reasoning by allowing the model to first perceive overall distribution information before generating the image, and improve generation stability by increasing the inference steps. IGTR shows an approximately 20% improvement over the AR baseline, demonstrating the effectiveness of reasoning prompts in complex image generation tasks.

### v3

@@ 4.2 Ablation Study
4.2.1 Robustness of VF Prompt
To evaluate the robustness of the proposed VF prompt, we conduct experiments using different types of VF prompts. Specifically, we use two types of prompts mentioned in Sec. 3.2.3: class VF prompt and universal VF prompt. We evaluate the effectiveness and robustness of VF prompt by comparing the results with consistent VF prompts (i.e., the same VF prompt used during both training and inference) and inconsistent VF prompts (i.e., different VF prompts used during training and inference). In addition, we also combine two different prompts for training and inference, with each having a 50% probability of being selected. We train the model with  [^1] M parameters for 300 epochs, which is the same setting as those in LlamaGen [37]. The baseline is LlamaGen [37], an AR image generation model based on the Llama architecture without our VF prompt.
@@ 4.2.1 Robustness of VF Prompt
Non VF prompt 8.69 124.44 0.79 0.47
@@
Full-view After Generation 8.70 122.79 0.79 0.47
VF prompt 6.07 174.48 0.82 0.44
@@
Table 4: Ablation study for different CFG settings.
The results are shown in Table 2, with the following observations: (1) Regardless of whether we use consistent or inconsistent VF prompts for training and inference, and regardless of which VF prompt we choose, our method consistently outperforms the baseline without VF prompt, achieving an approximate 20% improvement in FID scores. This demonstrates that our method exhibits good robustness. (2) The performance of the class VF prompt is better than that of the universal VF prompt. This is because focusing on a specific class is more informative than considering the overall distribution of the entire image set, as it more accurately reflects the representation of the specific class. However, the universal VF prompt is more generalizable as it is independent of the condition. (3) When the VF prompt is consistent between training and inference, the performance is better than when they are inconsistent. This is expected, as consistency between training and inference ensures that the model can better align the reasoning process across both phases, leading to more coherent and stable results. (4) The inferior performance of the mixed prompt is due to the introduction of information conflict and inconsistency. When using two different prompts, the model needs to balance both, which could lead to confusion in the generated content. Compared to a single prompt, the mixed prompt increases the complexity of reasoning, and the model may not have sufficient capacity to effectively integrate the information from both prompts, thus affecting performance.
@@ 4.2.2 Different Number of Prompt Tokens
We conduct an ablation study on the length of the intermediate VF prompt, using the universal VF prompt as the VF prompt, which samples a set of indices from the image tokenizer’s codebook based on a uniform distribution. To save computational resources, all experiments train the model with  [^1] M parameters for 50 epochs, as 50 epochs are sufficient to reflect the performance of different models. The results are shown in Fig. 4. A prompt length of 0 corresponds to the baseline, which represents the baseline without VF prompt. We can observe that when the VF prompt is short (8 tokens), the ability of VF prompt to handle complex image generation task is weaker because it contains fewer information, resulting in performance similar to the baseline. Furthermore, as the length of the prompt increases, the FID performance improves gradually. However, the performance gains are not uncapped, which can be found that when the prompt length reaches near 256, the gains begin to slow down. This phenomenon is reasonable as the gains from using representative information will reach an upper limit. Therefore, in subsequent experiments, considering both performance and efficiency, the length of VF prompt is set to 256.
@@
VF prompt-B (CFG=1.75)  [^1] M 4.39 222.80 0.85 0.46
VF prompt-L (CFG=1.75) 343M 3.23 279.07 0.84 0.52
VF prompt-XL (CFG=1.75) 775M 2.85 289.75 0.83 0.54
VF prompt-XXL (CFG=1.75) 1.4B 2.76 275.80 0.82 0.55
@@
Figure 5: Sample images randomly generated by VF prompt, trained on ImageNet.
Figure 6: Scaling behavior of VF prompt. (a) Training loss of models with different sizes. (b) FID scores of models with different sizes during inference.
Figure 7: Visualization of samples generated by VF prompt based on different model sizes.
@@ 4.2.3 Blank Prompt
Another way to think about our VF prompt is that the proposed image-related VF prompt enables the model to spend more computational resources (i.e., intermediate tokens). To isolate the impact of variable computation from the reasoning process itself, we conduct a variant where we replace the VF prompt with multiple blank tokens of equal length, which are tokens extracted from a completely black image with no representation information. The experiments train the model for 50 epochs. As shown in ‘Blank Prompt’ of Table 3, the experimental result indicates that this variant performs similarly to the baseline without our VF prompt. This demonstrates that variable computation itself is not the reason for the success of VF prompt, and proves the effectiveness of the designed VF prompt, which incorporates full-view image information.
4.2.4 Order of Full-view and Generation
A potential benefit of VF prompt is that it allows the model to better access relevant knowledge acquired during pretraining. To investigate this, we test an alternative configuration where the VF prompt is given only after the image generation, isolating whether the model depends on the prior VF prompt to generate the image. The experiments train the model for 50 epochs. As shown in the ‘Full-view After Generation’ row in Table 3, we find that the performance of this variant is comparable to the baseline, suggesting that the sequential full-view itself is valuable in the generation process of the model, rather than merely activating known knowledge.
@@ 4.2.5 Effect of Classifier-Free Guidance (CFG)
We conduct ablation experiments under different Classifier-Free Guidance (CFG) settings based on Universal VF prompt, as shown in Table 4. It can be observed that as CFG scale increases, the model’s performance gradually improves, with the best FID achieved when CFG scale is set to 1.75. However, further increasing CFG scale leads to a deterioration in FID. Additionally, the increase in CFG scale results in a trade-off between diversity and fidelity, with higher accuracy and lower recall. Therefore, in subsequent experiments, we set the CFG scale to 1.75.
@@ 4.3 Main Results
In Table 5, we compare with popular image generation models, including GAN [2, 13, 32], Diffusion models [6, 12, 28, 24], masked AR [4, 19, 20] and AR models [9, 44, 17, 37]. All our experiments are based on the universal VF prompt, as it is more general and easier to obtain compared to the class VF prompt due to its independence from the condition.
The results show that our models outperform all previous AR methods at different levels of model parameters. For example, compared to Llama-B, Llama-L Llama-XL and Llama-XL with the same parameters, VF prompt-B, VF prompt-L VF prompt-XL and VF prompt-XL improves FID by 20%, 15%, 16% and 11%, and IS by 15%, 12%, 28% and 9%, respectively. This indicates that our VF prompt of introducing VF prompts effectively helps the AR model in learning and generation by enhancing contextual reasoning of generation and improving stability during the generation process through increased inference steps . It is important to note that since LlamaGen does not release the results at 300 epoches for the L and XL models, for a fair comparison, we only present the results for our method at 50 epochs on the XL and XXL models, which have not yet reached the performance limit of our method, leaving room for further improvement.
@@
As shown in Fig. 5, we visualize some sample images randomly generated by VF prompt, trained on ImageNet. This demonstrates that VF prompt is capable of generating high-quality samples with both high fidelity and diversity. More visualizations are provided in the appendix.
@@
We investigate the scaling behavior of VF prompt. As shown in Fig. 6, we show the training loss and FID score based on models of different sizes. It can be observed that VF prompt demonstrates good scalability, with lower training loss and better FID scores as the model size increases. This is because we did not modify the formula or structure of the AR model itself, preserving the integrity of the AR framework, and therefore inheriting the scalability of AR methods. Additionally, we visualize the generation results of the same class under different model sizes, as shown in Fig. 7. It can be observed that as the model size increases, the generated results for the same class become more realistic and detailed, with finer textures and better overall consistency.
Figure 8: Training loss of the same model with and without VF prompt.
Convergence
As the loss curve shown in Fig. 8, it can be observed that the same model architecture after the same training step has a significantly smaller loss for the method using VF prompt, which confirms the conclusions of our theoretical analysis in Sec 3.2.2.
@@ 5 Conclusion
In this work, inspired by the prompt engineering from NLP of AR model, we propose the Vision Full-view (VF) prompt, to simulate the human process of creating images by guiding the model to first perceive vision full-view information before generating the image. This approach enhances contextual reasoning by providing a more comprehensive understanding of the overall image information. VF prompt also improves generation stability by increasing the inference steps. VF prompt shows an approximately 20% improvement over the AR baseline, demonstrating the effectiveness of VF prompts in complex image generation tasks.